OpenStudy (anonymous):

Solve the equation. 2( x + 4) = 30

OpenStudy (nenadmatematika):

x+4=15 so x=11

OpenStudy (lgbasallote):

distibute...occombine like terms...divide

OpenStudy (hoblos):

x+4=15 x=11

OpenStudy (anonymous):

thanks..happy valentines day:)

OpenStudy (lgbasallote):

im sorry for the typos...too lazy to edit them -___-

OpenStudy (anonymous):

very simple 2x+8=30 2x=30-8 2x=22 x=22/2 x=11